# Stability and Generalization of Quantum Neural Networks

### v1

@@ Abstract
Quantum neural networks (QNNs) play an important role as an emerging technology in the rapidly growing field of quantum machine learning. While their empirical success is evident, the theoretical explorations of QNNs, particularly their generalization properties, are less developed and primarily focus on the uniform convergence approach. In this paper, we exploit an advanced tool in statistical learning theory, i.e., algorithmic stability, to study the generalization of QNNs. We first establish high-probability generalization bounds for QNNs via uniform stability. Our bounds shed light on the key factors influencing the generalization performance of QNNs and provide practical insights into both the design and training processes. We next explore the generalization of QNNs on near-term noisy intermediate-scale quantum (NISQ) devices, highlighting the potential benefits of quantum noise. Moreover, we argue that previous analysis characterizes worst-case generalization guarantees, and we establish a refined optimization-dependent generalization bound for QNNs via on-average stability. Numerical experiments on various real-world datasets support our theoretical findings.
@@ 1 Introduction
In statistical learning theory, a variety of techniques for generalization analysis are known. One of the most popular approach is via the uniform convergence analysis, which studies the uniform generalization gaps in a hypothesis space. This approach typically uses complexity measures such as VC-dimension [22], covering numbers [23], Rademacher complexity [24] to develop capacity-dependent bounds. To the best of our knowledge, essentially all generalization bounds derived for QNNs so far are of the uniform kind, which, however, are not sufficient to explain the generalization behavior of current-scale QNNs [25]. Impressive alternatives have been proposed, including sample compression [26], PAC-Bayes [27], and algorithmic stability [28]. In particular, the algorithmic stability tools have advantages in some aspects, such as dimension-independent and adaptability for broad learning paradigms. Therefore, it is natural to explore the generalization of QNNs through the lens of algorithmic stability.
@@
  • We conduct numerical experiments on real-world datasets, following the seminal QML benchmark study [30]. The empirical results verify our theoretical findings.
@@ 2 Related Work
Generalization analysis of QNNs. The theory of generalization for QNNs is less developed and primarily focuses on the uniform convergence approach. Recent studies on the generalization of QNNs typically employ complexity measures such as pseudo-dimension [16], effective dimension [31], VC-dimension [21], Rademacher complexity [32, 18, 33], covering number [19, 20], and all their uniform relatives. Of the most interest to ours is [19], where the authors provide generalization bounds for QNNs based on covering number, all derived with the same setting as ours. Later, [20] uses quantum channels to derive more general results. Their generalization bounds exhibit a sublinear dependence on the number of trainable quantum gates. Unfortunately, this limits their results in providing meaningful guarantees for over-parameterized QNNs. More recently, it has been argued in [25] that the traditional measures of model complexity are not sufficient to explain the generalization behavior of current-scale QNNs. Their empirical findings highlight the need to shift perspective toward non-uniform generalization measures in QML. Therefore, it is natural to exploit algorithmic stability [28], which additionally enjoys desirable properties on flexibility (dimension-independence) and adaptivity (suiting for diverse learning scenarios).
@@ 3.2 Stability and Generalization Analysis of Randomized Algorithm
Algorithmic Stability. Algorithmic stability plays an important role in statistical learning, which measures the sensitivity of an algorithm to the perturbation of training sets. Our analysis relies on two widely used stability measures, namely uniform stability [28, 35] and on-average stability [37]. Below, we recall these notions.
@@
In the context of QNNs, the optimization typically employs SGD algorithm. For a given training dataset $S\in\mathcal{Z}^{m}$, the objective function to be minimized can be expressed as
@@ 5 Experimental Evaluation
Experiments Setup. Following the seminal QML benchmark study [30], we consider three datasets: Breast Cancer [61], MNIST [62], and Fashion MNIST [63]. We conduct the binary classification tasks on diagnosis B/M, digit 0/1, and class T-shirt/Trouser for the three datasets. The implementation of QNN is as follows. The classical information $\bm{x}$ is encoded into a quantum state $\rho(\bm{x})$ via angle encoding. We adopt the most widely used hardware-efficient ansatz such that the construction of $U(\bm{\theta})$ follows a layerwise structure using single-qubit Pauli rotation gates and two-qubit CNOT gates. We empirically estimate the generalization gap by calculating the absolute difference between the training and test errors. Each experiment setting is repeated 50 times to obtain statistical results.
Number of Trainable Quantum Gates. We investigate the impact of the number of trainable quantum gates $K$ on generalization gap by varying the QNN layer numbers $L=\{2,4,6,8,10\}$. The step size is set to $\eta=0.01$. In Figure 1, we observe that as $L$ increases, i.e., $K$ increases, the generalization gap also increases. This observation is consistent with our generalization bounds in Theorem 4.5 regarding the impact of $K$ on the generalization gap.
@@
Figure 1: Generalization gap for varying numbers of trainable quantum gates: (a) MNIST, (b) Fashion MNIST, and (c) Breast Cancer.
Figure 2: Generalization gap for varying step sizes: (a) MNIST, (b) Fashion MNIST, and (c) Breast Cancer.
Figure 3: Generalization gap with varying noise levels: (a) MNIST, (b) Fashion MNIST, and (c) Breast Cancer.
@@ 7 Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62102388), and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302901).

### v2

@@ Abstract
Quantum neural networks (QNNs) play an important role as an emerging technology in the rapidly growing field of quantum machine learning. While their empirical success is evident, the theoretical explorations of QNNs, particularly their generalization properties, are less developed and primarily focus on the uniform convergence approach. In this paper, we exploit an advanced tool in classical learning theory, i.e., algorithmic stability, to study the generalization of QNNs. We first establish high-probability generalization bounds for QNNs via uniform stability. Our bounds shed light on the key factors influencing the generalization performance of QNNs and provide practical insights into both the design and training processes. We next explore the generalization of QNNs on near-term noisy intermediate-scale quantum (NISQ) devices, highlighting the potential benefits of quantum noise. Moreover, we argue that our previous analysis characterizes worst-case generalization guarantees, and we establish a refined optimization-dependent generalization bound for QNNs via on-average stability. Numerical experiments on various real-world datasets support our theoretical findings.
@@ 1 Introduction
In classical learning theory, a variety of techniques for generalization analysis are known. One of the most popular approach is via the uniform convergence analysis, which studies the uniform generalization gaps in a hypothesis space. This approach typically uses complexity measures such as VC-dimension [22], covering numbers [23], Rademacher complexity [24] to develop capacity-dependent bounds. To the best of our knowledge, essentially all generalization bounds derived for QNNs so far are of the uniform kind, which, however, are not sufficient to explain the generalization behavior of current-scale QNNs [25]. Impressive alternatives have been proposed, including sample compression [26], PAC-Bayes [27], and algorithmic stability [28]. In particular, the algorithmic stability tools have advantages in some aspects, such as dimension-independent and adaptability for broad learning paradigms. Therefore, it is natural to explore the generalization of QNNs through the lens of algorithmic stability.
@@
  • We conduct numerical experiments on real-world datasets, and the empirical results indeed support our theoretical findings.
@@ 2 Related Work
Generalization analysis of QNNs. The theory of generalization for QNNs is less developed and primarily focuses on the uniform convergence approach. Recent studies on the generalization of QNNs typically employ complexity measures such as pseudo-dimension [16], effective dimension [30], VC-dimension [21], Rademacher complexity [31, 18, 32], covering number [19, 20], and all their uniform relatives. Of the most interest to ours is [19], where the authors provide generalization bounds for QNNs based on covering number, all derived with the same setting as ours. Later, [20] uses quantum channels to derive more general results. Their generalization bounds exhibit a sublinear dependence on the number of trainable quantum gates. Unfortunately, this limits their results in providing meaningful guarantees for over-parameterized QNNs. More recently, it has been argued in [25] that the traditional measures of model complexity are not sufficient to explain the generalization behavior of current-scale QNNs. Their empirical findings highlight the need to shift perspective toward non-uniform generalization measures in QML. Therefore, it is natural to leverage the concept of algorithmic stability [28], which additionally enjoys desirable properties on flexibility (dimension-independence) and adaptivity (suiting for diverse learning scenarios). While completing our work, we became aware of an anonymous work on OpenReview that derives a generalization bound for data re-uploading QNNs via uniform stability. We develop our approach independently of them, which differs in multiple key aspects and thus obtains more general results. Unlike their focus on constant step sizes, we also analyze the practically relevant scenarios of decaying step sizes. We further explore the generalization of QNNs in the noise scenario, highlighting the potential benefits of quantum noise. Moreover, we introduce another concept of on-average stability to provide the refined optimization-generalization bound.
@@ 3.2 Stability and Generalization Analysis of Randomized Algorithm
Algorithmic Stability. Algorithmic stability plays an important role in classical learning theory, which measures the sensitivity of an algorithm to the perturbation of training sets. Our analysis relies on two widely used stability measures, namely uniform stability [28, 34] and on-average stability [36]. Below, we recall these notions.
@@
Typically, SGD is employed as the classical optimizer in QNNs. Given a training dataset $S\in\mathcal{Z}^{m}$, the QNN minimizes the following objective function,
@@ 5 Experimental Evaluation
Experiments Setup. We consider two real-world datasets: MNIST [60] and Fashion MNIST [61], which have also been explored in the anonymous work as well as in other QML studies [62, 63, 64]. For these datasets, we conduct binary classification tasks on the digit 0/1 and the T-shirt/Trouser. The implementation of QNN is as follows. The classical information $\bm{x}$ is encoded into a quantum state $\rho(\bm{x})$ via angle encoding. We adopt the most widely used hardware-efficient ansatz such that the construction of $U(\bm{\theta})$ follows a layerwise structure using single-qubit Pauli rotation gates and two-qubit CNOT gates. We empirically estimate the generalization gap by calculating the absolute difference between the training and test errors. Each experiment setting is repeated 50 times to obtain statistical results.
Number of Trainable Quantum Gates. We investigate the impact of the number of trainable quantum gates $K$ on generalization gap by varying the QNN layer numbers $L=\{4,6,8,10,12\}$. The step size is set to $\eta=0.01$. In Figure 1, we observe that as $L$ increases, i.e., $K$ increases, the generalization gap also increases. This observation is consistent with our generalization bounds in Theorem 4.5 regarding the impact of $K$ on the generalization gap.
@@
Figure 1: Generalization gap for varying numbers of trainable quantum gates: (a) MNIST, (b) Fashion MNIST.
Figure 2: Generalization gap for varying step sizes: (a) MNIST, (b) Fashion MNIST.
Figure 3: Generalization gap with varying noise levels: (a) MNIST, (b) Fashion MNIST.
@@ 7 Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62102388), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302901), and the USTC Kunpeng & Ascend Center of Excellence.